OpenStudy (anonymous):

a)Find the solutions rounded in two decimal:{y=x^(-4/3),y=e^x b) A right triangle has an area of 21 square inches. The square of the hypotenuse is 85. Find the lengths of the legs of the triangle. Round your answer to the nearest inch.

OpenStudy (anonymous):

y=x^(-4/3),y=e^x common unknown is y and y=y x^(-4/3)=e^x

OpenStudy (anonymous):

1 equation and 1 unknown (x) solve for x